OpenStudy (marcelie):

Can someone please help me solve this ? For the following exercises, evaluate the function f at the indicated values f (−3), f (2), f (−a), − f (a), f (a + h). 27. f (x) = 2x − 5

OpenStudy (imtiaz7):

Okay; here we goo: f (x) = 2x − 5 For: 1) f(-3)= 2(-3)-5 =-6-5 =-11

OpenStudy (imtiaz7):

Follow above pattern to solve other questions, do you understand @marcelie

OpenStudy (marcelie):

Thank you !